# Towards Continual Expansion of Data Coverage: Automatic Text-guided Edge-case Synthesis

### v1

@@ Abstract
The performance of deep neural networks is strongly influenced by the quality of their training data. However, mitigating dataset bias by manually curating challenging edge cases remains a major bottleneck. To address this, we propose an automated pipeline for text-guided edge-case synthesis. Our approach employs a Large Language Model, fine-tuned via preference learning, to rephrase image captions into diverse textual prompts that steer a Text-to-Image model toward generating difficult visual scenarios. Evaluated on the FishEye8K object detection benchmark, our method achieves superior robustness, surpassing both naive augmentation and manually engineered prompts. This work establishes a scalable framework that shifts data curation from manual effort to automated, targeted synthesis, offering a promising direction for developing more reliable and continuously improving AI systems. Code is available at https://github.com/gokyeongryeol/ATES.
@@ 1 Introduction
These methods help mitigate deficiencies in existing datasets. However, they remain largely remedial in nature. A growing trend instead focuses on synthesizing new data to capture model weaknesses and underrepresented edge-case scenarios. Traditionally, identifying such edge-cases has been a labor-intensive process. It typically requires domain experts to manually analyze model failure modes, scrutinizing confusion matrices, class-wise performance metrics, and specific false positives and negatives, to guide subsequent data collection efforts [15]. This reliance on manual intervention renders the process not only time-consuming and expertise-dependent but also difficult to automate, scale, and reproduce. Please refer to Appendix A for a more comprehensive review of related studies.
@@ 2 Methodology
We aim to automatically synthesize high-value, edge-case data, thereby progressively expanding the coverage and robustness of a training dataset. The core of our methodology is a self-improving feedback loop wherein a LLM is fine-tuned to rephrase image captions, guiding a generative model toward producing images that are challenging for a given discriminative model. This process leverages preference learning to align the LLM’s rephrasing strategy with a quantitative measure of edge-ness. The overall pipeline is illustrated and detailed in Appendix B.
2.1 Initial data preparation: captioning and semantic expansion
@@
Next, to explore the semantic space around this description, $c_{base}$ is fed into a rephrasing LLM, denoted as $\pi_{\theta}$, where $\theta$ are its parameters. The LLM is prompted to produce a set of $N$ diverse semantic variations, $C=\{c_{1},c_{2},...,c_{N}\}$, where each $c_{i}\sim\pi_{\theta}(\cdot|c_{base})$. These rephrased captions are designed to alter contextual details, object attributes, or narrative perspectives of the original scene, forming the foundation for generating a diverse range of visual outputs. While increasing $N$ enhances the diversity of the rephrased captions, it also incurs additional computational cost, as each caption must be synthesized and pseudo-labeled in the following steps. For practicality, we fix $N=5$ in all experiments.
@@ 2.2 Quantifying edge-ness for preference labeling
A higher value of $s_{i}$ is interpreted as an indicator of edge-ness, as it signifies a greater discrepancy between the discriminative model’s predictions and the pseudo annotations.
2.3 Preference learning for edge-case-aware rephrasing
The constructed preference dataset $D_{pref}$ is used to fine-tune the rephrasing LLM $\pi_{\theta}$, aligning its behavior toward generating captions that are more likely to produce high edge-ness images. We employ Direct Preference Optimization (DPO) [29], a widely-used and effective algorithm for preference tuning. DPO directly optimizes the language model to satisfy the given preferences using the following loss function:
$\mathcal{L}_{DPO}(\pi_{\theta};\pi_{ref})=\mathbb{E}_{(c_{base},c_{w},c_{l})\sim D_{pref}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(c_{w}|c_{base})}{\pi_{ref}(c_{w}|c_{base})}-\beta\log\frac{\pi_{\theta}(c_{l}|c_{base})}{\pi_{ref}(c_{l}|c_{base})}\right)\right]$ (2)
where $\pi_{ref}$ is a reference copy of the initial LLM, $\sigma$ is the sigmoid function, and $\beta$ is a hyperparameter controlling the strength of the preference.
Once the LLM is aligned, the newly fine-tuned model $\pi_{\theta^{\prime}}$ is used to generate a set of edge-case-aware captions. These captions are used to synthesize new images, which, along with their pseudo annotations, are used to augment the original training set: $D_{aug}=D\cup\{(x_{new}^{\prime},y_{new}^{\prime})\}$. Notably, this entire pipeline can be executed iteratively. By retraining the discriminative model $M_{\phi}$ on $D_{aug}$ and using the updated model as the next-stage edge-ness scorer, the system can progressively discover and synthesize increasingly complex edge-cases, continually expanding the training data’s coverage.
@@ 3 Experiment
We aim to validate that our proposed pipeline effectively generates edge-case data that addresses the blind spots of a given discriminative model. We primarily compare the efficacy of different prompt generation strategies for augmenting a training dataset via the T2I model. Due to space constraint, details on model architectures and training configurations are provided in Appendix C.
3.1 Experimental setup
Comparison groups.
  • naive: This baseline directly uses the original, unmodified captions from the training dataset to generate new images, varying only the random seed of the T2I model for diversity.
Dataset.
We conduct our experiments on the FishEye8K dataset [8], a benchmark adopted to the AI-City Challenge. This dataset, captured from 20 fisheye cameras, contains annotations for five object classes (Bus, Bike, Car, Pedestrian, Truck) across diverse urban environments. Its known scene biases make it an ideal testbed for evaluating edge-case generation. As detailed in Appendix D, we partition the official training set into two disjoint subsets: train-D, used for training the discriminative model, and train-R, used for the preference learning of the rephrasing LLM. Note that the LLM is restricted from accessing train-D during preference learning and is exposed to it only at inference time. This ensures that the augmented dataset is synthesized based on the LLM’s generalization capability rather than any potential data leakage. For completeness, the test set is held out and used only for final evaluation.
Evaluation metrics.
We evaluate the effectiveness of the prompt generation strategies using the mAP gain of the discriminative model after augmenting the dataset with synthesized samples. To specifically assess how well each method addresses the model’s blind spots, we introduce a custom metric, named mAP w/o TP, which is detailed in Appendix E. This metric computes mAP exclusively on ground-truth annotations that were missed by the un-augmented discriminative model. A higher mAP w/o TP score indicates that the synthetic data more effectively improves the model’s initial blind spots.
3.2 Quantitative results
Table 1: YOLOv11-small performance with different training dataset configurations. Each synthesized dataset, denoted by an iteration tag of data augmentation, is expanded to 5 times of the train-D.
@@ 3.2 Quantitative results
As summarized in Table 1, our proposed method consistently outperforms both baselines across all experimental settings. When augmenting the dataset once (v1), our method achieves the highest mAP and, more importantly, the highest mAP w/o TP. [^1] Note that the model trained on train-D + naive v0 serves as the reference for calculating the mAP w/o TP. Accordingly, we measure this metric exclusively for subsequent iterations of data augmentations. This result demonstrates that the data generated by our pipeline is more effective at correcting the baseline model’s specific weaknesses compared to both naively generated data and data from manually engineered prompts. In contrast, other baselines show diminishing returns, suggesting that simply increasing data volume without targeting specific weaknesses is a less efficient strategy. For a class-wise and tag-wise analysis of the model performance to inspect the effect on dataset bias, please refer to Appendix F and G, respectively.
Furthermore, this performance advantage is maintained and even extended through a second iteration of data augmentation (v2). The consistent improvement of the proposed method highlights the efficacy of the self-improving loop, which progressively discovers and mitigates model blind spots.
Finally, we conducted transferability analysis introduced in recent data-centric approaches [20, 12, 7]. In Table 5, we reported the performance of YOLOv11-[medium, large] trained on the dataset curated with YOLOv11-small, where the dominance is maintained. For details, please refer to Appendix H.
3.3 Qualitative analysis
To understand the underlying properties of the generated data, we visualize the image embeddings obtained by CLIP [28] of each generation strategy using UMAP [21] in Figure 5. Compared to the naive, the manual exhibits a slightly broader spread, while the automatic not only achieves a greater advantage in terms of spread, but also shifts the mode toward a sparse region of real data embeddings.
@@ 3.3 Qualitative analysis
A closer inspection of the prompts and images of high-density regions reveals that our preference-tuned LLM learns a sophisticated rephrasing strategy. Compared to the factual source captions, its outputs tend to: (1) employ more descriptive, literary language; (2) emphasize actions and dynamic situations over static object descriptions; and (3) transform objective metadata (e.g., time, weather) into evocative descriptions of mood and atmosphere. This contrasts sharply with the manual baseline, where manual prompts are narrowly focused on object attributes. We interpret this as an automated strategy to enrich background context and scene diversity, which directly counteracts the inherent biases of the FishEye8K dataset. Please refer to Appendix I for more structured analysis.
@@ 4 Conclusion
Appendix A Related work
Synthetic data from generative models.
Automated synthesis of challenging data.
A significant body of research has focused on automatically generating challenging or out-of-distribution (OOD) data to improve model generalization. Early approaches, such as DREAM-OOD [5], learn a compact, text-conditioned latent space from in-distribution (ID) data. Outliers are then synthesized by sampling new embeddings from the low-likelihood regions of this learned manifold and subsequently decoding them into photorealistic images. This method effectively targets the boundaries of the known data distribution to produce novel samples.
Data-centric and prompt-based generation.
Appendix B Pipeline description
Figure 1: Training pipeline of the rephrasing LLM. Best viewed in color.
Training phase.
This diagram illustrates the process of preference-tuning the rephrasing LLM to understand and generate edge-case-aware captions. The process begins with an image sampled from the train-R split. An initial base caption is generated and subsequently rephrased into $N$ semantic variants by the LLM. Each of these $N$ captions is fed to a T2I model to synthesize a corresponding image, and the pseudo-labeler estimates the pseudo-annotations. A pre-trained discriminative model then evaluates all $N$ (synthetic image, and pseudo-annotations) pairs to quantify their "edge-ness," which is measured by the task loss. The captions that produced the maximum-loss (preferred) and minimum-loss (un-preferred) images are collected as a preference pair. This process is repeated to construct a preference dataset, which is then used to fine-tune the LLM’s Parameter-Efficient Fine-Tuning (PEFT) module via preference learning. This cyclical process teaches the LLM to generate prompts that are more likely to produce challenging images for the discriminative model.
@@ Appendix B Pipeline description
Appendix C Model architectures and training configurations
@@
All training was performed by the default configurations provided by each library, unless otherwise specified. We adopt most settings from prior work [15], and release all configuration files, bash scripts, and step-by-step guides for reproducing the experiments at the https://github.com/gokyeongryeol/ATES.
@@
Appendix D Analysis of bias in data splits

### v2

@@ Abstract
The performance of deep neural networks is strongly influenced by the quality of their training data. However, mitigating dataset bias by manually curating challenging edge cases remains a major bottleneck. To address this, we propose an automated pipeline for text-guided edge-case synthesis. Our approach employs a Large Language Model, fine-tuned via preference learning, to rephrase image captions into diverse textual prompts that steer a Text-to-Image model toward generating difficult visual scenarios. Evaluated on the FishEye8K object detection benchmark, our method achieves superior robustness, surpassing both naive augmentation and manually engineered prompts. This work establishes a scalable framework that shifts data curation from manual effort to automated, targeted synthesis, offering a promising direction for developing more reliable and continuously improving AI systems. Code is available at github repository.
@@ 1 Introduction
real naive manual automatic (ours)
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Comparison of synthetic data generation strategies. Each row displays examples generated from the same real image base caption using three different approaches: naive, manual, and automatic (ours). This row-wise arrangement isolates the effect of the generation strategy on the resulting visual output and its associated labels. Note that pseudo annotations are marked with different colors indicating different classes (Bus, Bike, Car, Pedestrian, Truck). Best viewed in color.
@@
These methods help mitigate deficiencies in existing datasets. However, they remain largely remedial in nature. A growing trend instead focuses on synthesizing new data to capture model weaknesses and underrepresented edge-case scenarios. Traditionally, identifying such edge-cases has been a labor-intensive process. It typically requires domain experts to manually analyze model failure modes, scrutinizing confusion matrices, class-wise performance metrics, and specific false positives and negatives, to guide subsequent data collection [15]. This manual intervention renders the process not only time-consuming and expertise-dependent but also difficult to automate, scale, and reproduce.
@@
2 Related work
Figure 2: Training pipeline of the rephrasing LLM to understand and generate edge-case-aware captions. This cyclical process teaches the LLM to generate prompts that are more likely to produce challenging images for the discriminative model. Best viewed in color.
Synthetic data from generative models.
Automated synthesis of challenging data.
A significant body of research has focused on automatically generating challenging or out-of-distribution (OOD) data to improve model generalization. Early approaches, such as DREAM-OOD [5], learn a compact, text-conditioned latent space from in-distribution (ID) data. Outliers are then synthesized by sampling new embeddings from the low-likelihood regions of this learned manifold and then decoding them into photorealistic images. This method targets the boundaries of the known data distribution to produce novel samples.
Data-centric and prompt-based generation.
@@ 3 Methodology
The overall pipeline is illustrated in Figure 2. We aim to automatically synthesize high-value, edge-case data, thereby progressively expanding the coverage and robustness of a training dataset. The core of our methodology is a self-improving feedback loop wherein a LLM is fine-tuned to rephrase image captions, guiding a generative model toward producing images that are challenging for a given discriminative model. This process leverages preference learning to align the LLM’s rephrasing strategy with a quantitative measure of edge-ness.
3.1 Initial data preparation
@@
Next, to explore the semantic space around this description, $c_{base}$ is fed into a rephrasing LLM, denoted as $\pi_{\theta}$, where $\theta$ are its parameters. The LLM is prompted to produce a set of $N$ diverse semantic variations, $C=\{c_{1},c_{2},...,c_{N}\}$, where each $c_{i}\sim\pi_{\theta}(\cdot|c_{base})$. These rephrased captions are designed to alter contextual details, object attributes, or narrative perspectives of the original scene, forming the foundation for generating a diverse range of visual outputs. While increasing $N$ enhances the diversity of the rephrased captions, it also incurs additional computational cost, as each caption must be synthesized and pseudo-labeled in the following steps. For practicality, we fix $N=5$.
@@ 3.2 Quantifying edge-ness for preference labeling
A higher value of $s_{i}$ is interpreted as an indicator of edge-ness, as it signifies a greater discrepancy between the model’s predictions and the pseudo annotations.
3.3 Edge-case-aware rephrasing
To align the LLM with the objective of generating challenging scenarios, we construct a preference dataset $D_{pref}$ consisting of pairs $(c_{w},c_{l})$, where $c_{w}$ denotes the caption that induced the maximum task loss (preferred) and $c_{l}$ denotes the caption that induced the minimum task loss (un-preferred) among the N variations generated for a given base caption. This dataset is used to fine-tune the rephrasing LLM $\pi_{\theta}$, effectively aligning its behavior toward generating captions that are more likely to produce high edge-ness images. For this preference-based fine-tuning, we adopt Direct Preference Optimization (DPO) [28], which optimizes the model to increase the relative log-probability of $c_{w}$ over $c_{l}$ using the following objective:
$\mathcal{L}_{DPO}$ $=\log\sigma\left(\beta\big(r_{\theta}(c_{w}\mid c_{base})-r_{\theta}(c_{l}\mid c_{base})\big)\right),$ (2)
$\text{where}\quad r_{\theta}(c\mid c_{base})=\log\frac{\pi_{\theta}(c\mid c_{base})}{\pi_{ref}(c\mid c_{base})}.$
Here, the reference model $\pi_{ref}$ is a frozen copy of the initial language model, $\sigma$ denotes the sigmoid function, and $\beta$ controls the strength of the preference signal.
Once the LLM is aligned, the newly fine-tuned model $\pi_{\theta^{\prime}}$ is used to generate a set of edge-case-aware captions. These captions are used to synthesize new images, which, along with their pseudo annotations, are used to augment the original training set: $D_{aug}=D\cup\{(x_{new}^{\prime},y_{new}^{\prime})\}$. Please refer to Appendix A for the data augmentation pipeline.
Notably, this entire pipeline can be executed iteratively. By retraining the discriminative model $M_{\phi}$ on $D_{aug}$ and using the updated model as the next-stage edge-ness scorer, the system can progressively discover and synthesize increasingly complex edge-cases, continually expanding the data coverage.
@@ 4 Experiment
Table 1: YOLOv11-small performance under iterative data augmentation strategies, comparing naive, manual, and automatic pipelines, with evaluation using mAP and mAP w/o TP.
@@
Table 2: Performance comparison for cross-scale transferability of synthetic data (small $\rightarrow$ medium/large).
(a) YOLOv11-medium performance.
Train dataset Count mAP mAP w/o TP
train-D 3,187 0.366 $\pm$ 0.004 -
train-D + naive v0 19,122 0.426 $\pm$ 0.002 -
train-D + naive v0 + naive v1 35,057 0.429 $\pm$ 0.004 0.415 $\pm$ 0.005
train-D + naive v0 + manual v1 35,057 0.428 $\pm$ 0.005 0.414 $\pm$ 0.005
train-D + naive v0 + automatic v1 35,057 0.429 $\pm$ 0.003 0.416 $\pm$ 0.005
(b) YOLOv11-large performance.
Train dataset Count mAP mAP w/o TP
train-D 3,187 0.371 $\pm$ 0.001 -
train-D + naive v0 19,122 0.419 $\pm$ 0.001 -
train-D + naive v0 + naive v1 35,057 0.431 $\pm$ 0.003 0.414 $\pm$ 0.004
train-D + naive v0 + manual v1 35,057 0.429 $\pm$ 0.007 0.411 $\pm$ 0.008
train-D + naive v0 + automatic v1 35,057 0.432 $\pm$ 0.005 0.415 $\pm$ 0.006
We aim to validate whether our proposed pipeline effectively generates edge-case data that addresses the blind spots of a given discriminative model. We primarily compare the efficacy of different prompt generation strategies for augmenting a training dataset via the T2I model. Due to space constraint, details on model architectures and training configurations are provided in Appendix B.
4.1 Setup
Comparison groups.
  • naive: This baseline directly uses the base captions ($c_{base}$) generated by InternVL3-38B [35] from the training images. It generates new images without any further rephrasing or modification to these captions, varying only the random seed of the T2I model for diversity.
Dataset.
We conduct our experiments on the FishEye8K dataset [8], a benchmark adopted to the AI-City Challenge. The dataset consists of images from 20 fisheye cameras with annotations for five categories: Bus, Bike, Car, Pedestrian, and Truck. Given its inherent scene biases across diverse urban environments, FishEye8K serves as an optimal testbed for evaluating the effectiveness of our automated edge-case synthesis.
To rigorously assess model robustness, we partition the official training set into two disjoint subsets: train-D and train-R. To simulate a restricted training environment, train-D is used for training the discriminative model while train-R is used for the preference learning of the rephrasing LLM and contains a more diverse distribution. This setup allows the LLM to learn under-represented characteristics and generate corresponding synthetic data to fill the gaps in train-D. Please refer to Appendix C for more details.
To prevent potential data leakage, the LLM is restricted from accessing train-D during the preference learning phase and only interacts with it at inference time for data augmentation. This protocol ensures that the synthesized samples reflect the LLM’s generalization capability rather than simple memorization. For completeness, the test set is held out exclusively for final evaluation.
Evaluation metrics.
We evaluate the effectiveness of the prompt generation strategies using the mAP gain of the discriminative model after augmenting the dataset with synthesized samples. This evaluation protocol allows us to directly quantify how much each augmentation strategy contributes to improving detection performance beyond the original training data. To explicitly assess how well each method addresses the model’s blind spots, we introduce a custom metric, termed mAP w/o TP, described in Appendix D. In brief, this metric computes mAP only on ground-truth annotations that do not overlap with the True Positive predictions of the discriminative model trained on train-D and naive v0, thereby focusing exclusively on previously misrecognized or challenging instances. A higher mAP w/o TP therefore indicates that the synthetic data more effectively improves the model’s initial blind spots, rather than merely reinforcing patterns that the model has already learned.
4.2 Visual diversity
Figure 1 provides a controlled comparison of how different prompt generation strategies influence synthetic visual outputs. By utilizing the same base caption for each row, we can directly observe the impact of rephrasing on synthesis.
While the automatic approach does not necessarily generate a higher number of object instances than the manual variant, it exhibits significantly greater background diversity, encompassing a wider range of lighting conditions, camera viewpoints, and scene contexts. This suggests that the preference-tuned LLM successfully explores a broader distribution of plausible scenarios rather than sticking to the main modes of the training distribution. In contrast, the manual baseline often produces repetitive scenes with limited background variation. Consequently, the automatic method is more effective at generating novel samples that cover underrepresented regions of the data distribution, thereby helping to alleviate dataset bias and improve model robustness
4.3 Model performance
@@
As summarized in Table 1, our proposed method consistently outperforms both baselines across all experimental settings. To establish a competitive baseline, we define v0 as the initial iteration of data augmentation, where the train-D is expanded five-fold using the naive strategy. This naive v0 dataset serves as the common starting point for all subsequent augmentation experiments. When augmenting the dataset further in the next iteration (v1), our method achieves the highest mAP and, more importantly, the highest mAP w/o TP. [^1] Note that the model trained on train-D + naive v0 serves as the reference for calculating the mAP w/o TP. Accordingly, we measure this metric exclusively for subsequent iterations of data augmentations.
This result demonstrates that the data generated by our pipeline is more effective at correcting the baseline model’s specific weaknesses compared to both naively generated data and data from manually engineered prompts. In contrast, other baselines show diminishing returns, suggesting that simply increasing data volume without targeting specific weaknesses is a less efficient strategy. For a class-wise and tag-wise analysis of the model performance to inspect the effect on dataset bias, please refer to Appendix E and F, respectively.
4.4 Transferability
To assess whether a dataset deemed effective for one model remains valid for another, a core concern in data-centric evaluation, we conducted transferability experiments. This approach aligns with recent findings [12, 7] that emphasize robust evaluation designs, acknowledging that instance informativeness may vary across model configurations and necessitating experimental methods to identify broadly transferable data. To examine the transferability of synthetic data across model scales, we trained YOLOv11-medium and YOLOv11-large using datasets generated by the proposed pipeline with YOLOv11-small. As summarized in Table 2(b), the automatic still achieves the highest mAP and mAP w/o TP.
Figure 3: Comparison of edge-ness statistics across different data augmentation strategies. The proposed automatic pipeline produces samples with the highest mean and median loss, indicating more challenging training examples for the discriminative model.
(a) naive
(b) manual
(c) automatic
Figure 4: Comparison of naive, manual, and automatic by UMAP of CLIP embeddings. Gray points represent real data, while the colored ones indicate synthetic data, with colors closer to yellow denoting higher density.
4.5 Edge-ness
We empirically validate the core mechanism of our pipeline: its ability to generate images with high edge-ness. We measure the training loss $\mathcal{L}_{task}$ of the discriminative model on data generated by each method. Following the standard training objective of our YOLOv11-small discriminator, $\mathcal{L}_{task}$ is specifically defined as the weighted sum of the bounding box regression loss, classification loss, and distribution focal loss. [14]. As shown in Figure 3, data from our proposed pipeline consistently induces the highest mean and median loss. This confirms that our preference learning framework successfully steers the LLM to generate captions that translate into verifiably more difficult examples for the discriminative model, thereby creating effective training signals for improving robustness.
4.6 Embedding
To better understand the underlying properties of the generated data, we visualize the image embeddings obtained by CLIP [27] for each generation strategy using UMAP [20], as shown in Figure 4. This visualization provides an intuitive view of how different strategies explore the feature space induced by a strong vision-language model. Compared to the naive approach, the manual strategy exhibits a slightly broader spread, indicating a modest increase in diversity. In contrast, the automatic method not only achieves a significantly greater spread, but also shifts its mode toward a sparser region of the real data embedding space. This shift suggests that the automatic strategy is more effective at generating novel and less frequently observed samples, potentially covering underrepresented regions of the data distribution. Such behavior is consistent with our objective of synthesizing edge-case examples that complement the existing training data.
4.7 Linguistic analysis
@@
A closer inspection of the prompts and images of high-density regions reveals that our preference-tuned LLM learns a sophisticated rephrasing strategy. Compared to the factual source captions, its outputs tend to: (1) employ more descriptive, literary language; (2) emphasize actions and dynamic situations over static object descriptions; and (3) transform objective metadata (e.g., time, weather) into evocative descriptions of mood and atmosphere. This contrasts sharply with the manual baseline, where manual prompts are narrowly focused on object attributes. We interpret this as an automated strategy to enrich background context and scene diversity, which directly counteracts the inherent biases of the FishEye8K dataset. Please refer to Appendix G for more structured analysis.
@@ 5 Conclusion
Appendix A Data augmentation pipeline
@@
Appendix B Model architectures and training configurations
@@
All training was performed by the default configurations provided by each library, unless otherwise specified. We adopt most settings from prior work [15], and release all configuration files, bash scripts, and step-by-step guides for reproducing the experiments at the github repository.
@@
Appendix C Analysis of bias in data splits